HUBBLE SPACE TELESCOPE DAILY REPORT #1872 PERIOD COVERED: 0000Z (UTC) 05/06/97 - 0000Z (UTC) 05/07/97 Daily Status Report as of 127/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of NIC/1/2/3 7043 (NICMOS Focus Monitor) The Near Infrared Cameras (Cameras One, Two and Three) were used to make observations of NGC-3603 in order to make focus calibration and monitoring measurements. This proposal is built to run NIC/1/2 focus sweeps on a weekly basis, and NIC/3 focus sweeps twice a week during SMOV. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of WF/PC-2 7553 (Pure Parallel Archive Survey II) WF/PC-2 parallel observations were obtained in the ultraviolet in search for interesting objects. This program is designed to find extragalactic sources with near ultraviolet excesses (3000 A0) outside the galactic plane and far ultraviolet sources (1600 A0) in the galactic plane. The procedure was executed as scheduled, and no anomalies were reported. 1.3 Completed Two Sets of WF/PC-2 6909 (WF/PC-2 Cycle 6 Calibration: Earth Flats The WF/PC-2 was used to make observations of the bright Earth to obtain streak flats which are used to construct high quality flat fields for the WF/PC-2 filter set. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed NIC/2 7154 (NICMOS Coronographic Hole Monitor) The Near Infrared Camera (Camera Two) was used to make observations of object NGC-3606 in order to monitor the effects of coronagrapghic hole. The purpose of this proposal is to take lamp flats with camera 2 (and cameras 1 and 3 occasionally) to monitor the location of the coronographic spot. The observations were completed as scheduled, and no anomalies were reported. 1.5 Completed Two Sets of WF/PC-2 6346 (Snapshot Survey of the Ultraluminous IRAS Galaxy Sample-Part 1 High) The WF/PC-2 was used to make snapshot survey observations of the ultraluminous IRAS galaxies IR20551-42 and NGWULIR143. These data will be searched for multiple cores, which imply a merger has occurred, and for starburst regions. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Two Sets of WF/PC-2/GHRS/1 6743 (HST Far UV Imaging and Spectra of Jupiter's Aurora) The WF/PC-2 and the High Resolution Spectrograph (Side One) were both used to make fat-ultraviolet observations of the auroral zones on the planet Jupiter. These observations will be compared with observations from the Galileo Probe take at about the same time. The observations were executed as scheduled, and no anomalies were noted. 1.7 Completed S/C/NIC/1/2/3 7037 (NICMOS Transfer Function Verification Test) A test was conducted to verify the basic operating characteristics of the NICMOS FPAs. This was accomplished by taking a series of dark frames at different detector bias voltages (0.0-0.6V in 0.1V steps). These were taken in parallel as the data was directly compared with ground test data. The observations were completed as scheduled, and no anomalies were noted. 1.8 Completed WF/PC-2 6632 (Morphologies of High Redshifted Radio Galaxies) The WF/PC-2 was used to make observations of the high redshifted radio galaxy GAL-2144+1 in order to study its morphology. This is part of a study of the continuum and Lyman-Alpha morphologies of 20 high redshifted radio galaxies which was begun in Cycle 5. The observations were completed as planned, and no problems were reported. 1.9 Completed WF/PC-2 6887 (A Photometric Search for Faint Companions to Nearby Stars) The WF/PC-2 was used to make a photometric search observations of the nearby star HD187642 in order to look for possible companions. The technique involves use of the Planetary Camera with wide band pass filters to photometrically measure for the presence of resolved dark companions. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 11 Successful : 11 Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled : 21 Successful : 21 2.3 Operations Notes: The battery #5 capacity check test was started at 126/1815Z. By 127/0919Z, the pressure was 410 psi and the voltage was 20.44. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science and SMOV operations. /DMH